# Latent-space variational data assimilation in two-dimensional turbulence

### v1

@@ 4.1 Accuracy of the estimated fields
To understand the cause of the high-wavenumber artefacts in state-space DA, it should be noted that the adjoint equations (3) are forced by $\mathcal{D}\mathcal{J}/\mathcal{D}\omega$, which is a series of singular impulses in space and time at each measurement location. These superpositions of delta functions are advected and diffused by (3), but signatures of this forcing are apparent in the visualisation of $\omega^{\dagger}(t=0)$ in figure 4(bi). As the Fourier transform of a delta function is a constant function of $k$, these singular impulses lead to high wavenumber spectral content of the state-space adjoint variables (figure 4(a)). Each variational DA iteration slightly perturbs the turbulent field estimate in a direction with high energy in the high wavenumbers. For latent-DA, the effective adjoint update can be visualised in state space by evaluating $\mathcal{F}_{\mathcal{D}}(\eta+\alpha\eta^{\dagger})-\mathcal{F}_{\mathcal{D}}(\eta)$. As shown in figure 4, the energy in the high wavenumbers of this effective latent adjoint direction is reduced to machine precision, which is consistent with the improved spectrum of $\omega_{0}^{*}$ for LatentDA.
4.2 Observability in latent space
@@
it is clear that the columns of the decoder Jacobian are the perturbation directions which, to linear order, approximately remain on the turbulent attractor. The linear transformation of adjoints from state to latent space (6) can now be understood as a projection onto these physically relevant perturbation directions. The associated linearised update in state space is given by,
@@
As such, the estimate remains physically relevant in state space throughout the latent variational DA method, to linear order. Physically, these improved perturbation directions are marked by an absence of the high-wavenumber artefacts discussed previously.
The second perspective considers how observable the reference initial turbulent field is from the measurements. As we will show here, the iterative gradient-based updates throughout the variational DA method are also expansions in some basis of adjoint fields. When assimilating data variationally, it is beneficial for the reference turbulent field $\omega^{R}_{0}$ to be well represented in this basis, such that the iterative updates can more effectively converge onto $\omega^{R}_{0}$. To begin, we define the deviation field $w=\omega-\omega^{R}$ which is governed by the linearised Navier-Stokes equations
@@
with the associated deviation streamfunction $\varphi=-\nabla^{2}w$ and the ground truth streamfunction $\psi^{R}=-\nabla^{2}\omega^{R}$. As shown in wang2022observable for turbulent channel flow, the DA cost function (2) can be written in terms of the measurement kernel $\phi(\bm{x}_{m})$ which extracts the measurement of interest at location $\bm{x}_{m}$
@@
$\left\langle w(t=t_{n}),\phi(\bm{x}_{m})\right\rangle=\left\langle\mathcal{L}w_{0},\phi(\bm{x}_{m})\right\rangle=\left\langle w_{0},\mathcal{L}^{\ddagger}\phi(\bm{x}_{m})\right\rangle=\left\langle w_{0},w^{\ddagger}(t=0;t=t_{n},\bm{x}_{m})\right\rangle,$
where $\mathcal{L}$ is the forward operator of the linearised Navier-Stokes equations (10) which advances $w_{0}$ to $w(t=t_{n})$. Note that $w^{\ddagger}\neq\omega^{\dagger}$, and an explicit relation between them will be given below. The associated adjoint operator $\mathcal{L}^{\ddagger}$ solves the linearised adjoint equations
$\frac{\partial w^{\ddagger}}{\partial\tau}+J(\psi^{R},w^{\ddagger})-\frac{1}{Re}\nabla^{2}w^{\ddagger}+\varphi^{\ddagger}=0\,,\quad\nabla^{2}\varphi^{\ddagger}-J(\omega^{R},w^{\ddagger})=0\,$, (12)
@@
enabling the explicit expression of the gradient,
@@
The relative error $\varepsilon$ of the reconstruction is reported in figure 5, for $Re=100$, where we compare SR-DA (green) and $\eta_{0}^{*}$ obtained with LatentDA (red). The first 500 adjoint POD modes were computed by the Arnoldi iteration. The reconstructions were also time-evolved over the DA time horizon, and the time-averaged relative errors (circles) were computed as a function of the number of adjoint POD modes used. The reference turbulent state can be reconstructed to ${\sim}5\%$ relative error in the latent adjoint basis with 500 adjoint POD modes, as opposed to ${\sim}50\%$ relative error in the state-space adjoint basis, demonstrating the improved observability of $\omega^{R}_{0}$ from the measurements in the latent space near optimality. Perhaps more notably, ${\sim}100$ modes in latent space can reconstruct the reference turbulent state with ${\sim}20\%$ error, while the same number of modes can only reconstruct the state with ${\sim}75\%$ error in state space. High wavenumber artefacts and signatures of the localized adjoint forcing are evident in the state-space adjoint basis reconstruction (figure 5(bi)). These small-scale fluctuations decay over the DA time horizon, explaining the improved time-averaged relative errors in state space. In contrast, the decoded reconstruction in latent space very closely resembles the reference turbulent state (panels bii-biii), and the error in the representation of the initial state and the time-averaged errors during the evolution are similar. These results demonstrate that the observability of the turbulence using limited measurements can be appreciably improved when the assimilation is performed in the right space, or coordinates.
5 Conclusion
@@
In state-space DA, the adjoint field is highly localised in the vicinity of the delta-function forcing, such that the spectrum of the adjoint field is broadband. The updates in latent space are more targetted and physically relevant, resulting in more accurate reconstruction of both the large and small scales, and a smaller departure from the truth over time. An improvement of an order of magnitude was achieved when assimilating data in latent space at $Re=40$ and 100, compared to the best approach considered in state space. A lower but nonetheless appreciable improvement was achieved at $Re=400$, where the increased spatiotemporal complexity of the flow requires a larger latent space dimensionality. The small scales of the estimated flow state are more dynamically relevant at all $Re$ considered.
@@
{bmhead}
[Declaration of interests] The authors report no conflict of interest.
{bmhead}
[Supplementary material] Supplementary material is available at …

### v2

@@ IV.1 Accuracy of the estimated fields
To understand the cause of the high-wavenumber artifacts in state-space DA, it should be noted that the adjoint equations (3) are forced by $\mathcal{D}\mathcal{J}/\mathcal{D}\omega$, which is a series of singular impulses in space and time at each measurement location. These superpositions of delta functions are advected and diffused by (3), but signatures of this forcing are very apparent in the visualization of $\omega^{\dagger}(t=0)$ at $Re=400$ in figures 4(bi,bii). As the Fourier transform of a delta function is a constant function of $k$, these singular impulses lead to high wavenumber spectral content of the state-space adjoint variables (figure 4a). Each variational DA iteration slightly perturbs the turbulent field estimate in a direction with high energy in the high wavenumbers. For latent-DA, the effective adjoint update can be visualized in state space by evaluating $\mathcal{F}_{\mathcal{D}}(\eta+\alpha\eta^{\dagger})-\mathcal{F}_{\mathcal{D}}(\eta)$. As shown in figure 4, the energy in the high wavenumbers of this effective latent adjoint direction is many orders of magnitude smaller than the state-space adjoint directions, which is consistent with the improved spectrum of $\omega_{0}^{*}$ for LatentDA. Adjoint directions are normalized before their spectra are computed to enable comparison between state space and latent space.
IV.2 Robustness to noisy measurements
Figure 5: Relative error of the estimated turbulent field at (left) the initial time $\varepsilon(t=0)$ and (right) the DA time horizon $\varepsilon(t=T)$ as a function of the noise amplitude $\sigma$, by (blue) InterpDA, (orange) SR, (green) SR-DA and (red) LatentDA at $Re=400$ (markers are slightly offset for clarity).
In practice, the measurement data available for assimilation are typically not perfect, but rather corrupted by some sensor error. We briefly investigate the effect of measurement noise on LatentDA as a test of the robustness of the proposed method. Similarly to Wang and Zaki [22] and Page [19], noise is added to the measurements of the reference trajectories
$m^{R}_{n,j}\leftarrow m^{R}_{n,j}+\Sigma_{n,j},$ (10)
where $m^{R}_{n}=\{m^{R}_{n,j}\}_{j=1}^{N_{x}N_{y}/M^{2}}$ are the measurements at each discrete time $t_{n}$, and the noise is normally distributed based on the local absolute vorticity $\Sigma_{n,j}\sim\mathcal{N}(0,\sigma|m^{R}_{n,j}|)$. We consider noise amplitudes of $\sigma\in\{0,0.05,0.1\}$, and the same noise is also present when initializing each variational procedure, whether by interpolation (InterpDA) or by a pre-trained SR network inference (SR, SR-DA and LatentDA).
The relative error of the estimated turbulent state at the initial time $t=0$ and at the DA time horizon $t=T$ are reported as a function of $\sigma$ in figure 5 for $Re=400$. All methods are similarly robust to the increasing noise at both times considered. As such, LatentDA maintains its reconstruction accuracy advantage over all state-space assimilation methods when measurements are noisy. This is a promising result in the context of scaling LatentDA to higher-$Re$, experimental measurements.
IV.3 Observability in latent space
@@
it is clear that the columns of the decoder Jacobian are the perturbation directions which, to linear order, approximately remain on the turbulent attractor. The linear transformation of adjoints from state to latent space (6) can now be understood as a projection onto these physically relevant perturbation directions. The associated linearized update in state space is given by,
@@
As such, the estimate remains physically relevant in state space throughout the latent variational DA method, to linear order. Physically, these improved perturbation directions are marked by an absence of the high-wavenumber artifacts discussed previously.
The second perspective considers how observable the reference initial turbulent field is from the measurements. As we will show here, the iterative gradient-based updates throughout the variational DA method are also expansions in some basis of adjoint fields. When assimilating data variationally, it is beneficial for the reference turbulent field $\omega^{R}_{0}$ to be well represented in this basis, such that the iterative updates can more effectively converge onto $\omega^{R}_{0}$. To begin, we define the deviation field $w=\omega-\omega^{R}$ which is governed by the linearized Navier-Stokes equations
@@
with the associated deviation streamfunction $\varphi=-\nabla^{2}w$ and the ground truth streamfunction $\psi^{R}=-\nabla^{2}\omega^{R}$. As shown in Wang et al. [23] for turbulent channel flow, the DA cost function (2) can be written in terms of the measurement kernel $\phi(\bm{x}_{m})$ which extracts the measurement of interest at location $\bm{x}_{m}$
@@
$\left\langle w(t=t_{n}),\phi(\bm{x}_{m})\right\rangle$ $=\left\langle\mathcal{L}w_{0},\phi(\bm{x}_{m})\right\rangle$
$=\left\langle w_{0},\mathcal{L}^{\ddagger}\phi(\bm{x}_{m})\right\rangle$
$=\left\langle w_{0},w^{\ddagger}(t=0;t=t_{n},\bm{x}_{m})\right\rangle,$
where $\mathcal{L}$ is the forward operator of the linearized Navier-Stokes equations (11) which advances $w_{0}$ to $w(t=t_{n})$. Note that $w^{\ddagger}\neq\omega^{\dagger}$, and an explicit relation between them will be given below. The associated adjoint operator $\mathcal{L}^{\ddagger}$ solves the linearized adjoint equations
$\frac{\partial w^{\ddagger}}{\partial\tau}+J(\psi^{R},w^{\ddagger})-\frac{1}{Re}\nabla^{2}w^{\ddagger}+\varphi^{\ddagger}$ $=0\,$, (13a)
$\nabla^{2}\varphi^{\ddagger}-J(\omega^{R},w^{\ddagger})$ $=0\,$, (13b)
@@
enabling the explicit expression of the gradient and the Hessian of the DA cost function,
@@
The relative error $\varepsilon$ of the reconstruction is reported in figure 6, for $Re=100$, where we compare SR-DA (green) and $\eta_{0}^{*}$ obtained with LatentDA (red). The first 500 adjoint POD modes were computed by the Arnoldi iteration. The reconstructions were also time-evolved over the DA time horizon, and the time-averaged relative errors (circles) were computed as a function of the number of adjoint POD modes used. The reference turbulent state can be reconstructed to ${\sim}5\%$ relative error in the latent adjoint basis with 500 adjoint POD modes, as opposed to ${\sim}50\%$ relative error in the state-space adjoint basis, demonstrating the improved observability of $\omega^{R}_{0}$ from the measurements in the latent space near optimality. Perhaps more notably, ${\sim}100$ modes in latent space can reconstruct the reference turbulent state with ${\sim}20\%$ error, while the same number of modes can only reconstruct the state with ${\sim}75\%$ error in state space. High wavenumber artifacts and signatures of the localized adjoint forcing are evident in the state-space adjoint basis reconstruction (figure 6bi). These small-scale fluctuations decay over the DA time horizon, explaining the improved time-averaged relative errors in state space. In contrast, the decoded reconstruction in latent space very closely resembles the reference turbulent state (panels bii-biii), and the error in the representation of the initial state and the time-averaged errors during the evolution are similar. These results demonstrate that the observability of the turbulence using limited measurements can be appreciably improved when the assimilation is performed in the right space, or coordinates.
V Conclusion
@@
In state-space DA, the adjoint field is highly localized in the vicinity of the delta-function forcing, such that the spectrum of the adjoint field is broadband. The updates in latent space are more targeted and physically relevant, resulting in more accurate reconstruction of both the large and small scales, and a smaller departure from the truth over time. An improvement of an order of magnitude was achieved when assimilating data in latent space at $Re=40$ and 100, compared to the best approach considered in state space. A lower but nonetheless appreciable improvement was achieved at $Re=400$, where the increased spatiotemporal complexity of the flow requires a larger latent space dimensionality. The small scales of the estimated flow state are more dynamically relevant at all $Re$ considered. LatentDA also exhibited a comparable robustness to noisy measurements as all state-space DA methods.
@@
Appendix A Continuous adjoint equations: vorticity-streamfunction formulation
The 2D velocity-vorticity Navier-Stokes equations (1) with some forcing $g$ can be written in terms of the vorticity and streamfunction $\psi$ as
$\frac{\partial\omega}{\partial t}-J(\psi,\omega)-\frac{1}{Re}\nabla^{2}\omega+g=0,$ (18)
where
$J(\psi,\omega)=\frac{\partial\psi}{\partial x}\frac{\partial\omega}{\partial y}-\frac{\partial\psi}{\partial y}\frac{\partial\omega}{\partial x}$ (19)
and $\psi$ is related to the velocity components via $u=\partial_{y}\psi,v=-\partial_{x}\psi$, such that $\psi$ is recovered from $\omega$ by solving the Poisson equation $\nabla^{2}\psi=-\omega$. The constrained optimization problem (2) can be cast as an unconstrained optimization problem by defining the Lagrangian
$\mathscr{L}=\mathcal{J}-\left(\omega^{\dagger},\frac{\partial\omega}{\partial t}-J(\psi,\omega)-\frac{1}{Re}\nabla^{2}\omega+g\right)-\left(\psi^{\dagger},\omega+\nabla^{2}\psi\right),$ (20)
where the inner product is defined by integrating over space $V$ and the observation time horizon $t\in[0,T]$
$\left(a,b\right)=\int_{0}^{T}\int_{V}ab\,\,dVdt.$ (21)
Taking variations of $\mathscr{L}$ with respect to the Lagrange multipliers $\omega^{\dagger}$ and $\psi^{\dagger}$ yields the Navier-Stokes equation for the estimated $\omega$ and the Poisson equation relating $\omega$ to $\psi$. To take variations of $\mathscr{L}$ with respect to $\omega$ and $\psi$, we proceed as follows. Assuming periodic boundary conditions, one can re-express terms such as $\left(\omega^{\dagger},J(\psi,\omega)\right)$ by integrating by parts in space
$\left(\omega^{\dagger},J(\psi,\omega)\right)$ $=\int_{0}^{T}\int_{V}\omega^{\dagger}\left[\frac{\partial\psi}{\partial x}\frac{\partial\omega}{\partial y}-\frac{\partial\psi}{\partial y}\frac{\partial\omega}{\partial x}\right]\,\,dVdt,$
$=\iint\left[\frac{\partial}{\partial x}\left(\omega^{\dagger}\frac{\partial\psi}{\partial y}\right)\omega-\frac{\partial}{\partial y}\left(\omega^{\dagger}\frac{\partial\psi}{\partial x}\right)\omega\right]dVdt,$
$=-\left(J(\psi,\omega^{\dagger}),\omega\right),$
where in the final equality, we assume commutativity of the partial derivatives. By the antisymmetry of $J$, the relation $\left(\omega^{\dagger},J(\psi,\omega)\right)=\left(J(\omega,\omega^{\dagger}),\psi\right)$ follows quickly. By isolating $\omega$ in the remaining terms, the Lagrangian (20) can be written as
$\mathscr{L}=\mathcal{J}$- $\omega^{\dagger}(T)\omega(T)+\omega^{\dagger}(0)\omega_{0}$
$+\left(\frac{\partial\omega^{\dagger}}{\partial t}-J(\psi,\omega^{\dagger})+\frac{1}{Re}\nabla^{2}\omega^{\dagger}-\psi^{\dagger},\omega\right)$
$-\left(\omega^{\dagger},g\right)-\left(\nabla^{2}\psi^{\dagger},\psi\right).$ (22)
Taking variations of $\mathscr{L}$ with respect to $\omega$ is now straightforward, and variations with respect to $\psi$ follow quickly from the antisymmetry of $J$. Setting $\mathcal{D}\mathscr{L}/\mathcal{D}\omega=0$ and $\mathcal{D}\mathscr{L}/\mathcal{D}\psi=0$ yields the equations for the adjoint fields $\omega^{\dagger}$ and $\psi^{\dagger}$
$\frac{\partial\omega^{\dagger}}{\partial\tau}+J(\psi,\omega^{\dagger})-\frac{1}{Re}\nabla^{2}\omega^{\dagger}+\psi^{\dagger}$ $=\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\omega},$ (23)
$\nabla^{2}\psi^{\dagger}-J(\omega,\omega^{\dagger})$ $=\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\psi},$ (24)
where $\tau=T-t$ is the backwards time. Then, the adjoint field at $t=0$ yields the total variation of the cost function with respect to the initial condition $\omega_{0}$,
$\frac{\mathscr{D}\mathcal{J}}{\mathscr{D}\omega_{0}}=\omega^{\dagger}(t=0).$ (25)
Appendix B IRMAE architecture
B.1 $Re=40$ & 100100
The IRMAE architecture used at $Re=40$ and $100$ is adopted from Cleary and Page [4]. Dense blocks (DB) [12] are groups of convolutional layers where the output of each layer is concatenated with its input. Here, each dense block is made up of three convolutional layers, each with 32 filters. We apply the ‘GELU’ activation function [11] and batch normalization [13] to the output of convolutional layers. The encoder architecture is given by the following sequence of operations:
$\omega$ $\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)\to\text{MP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{MP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{MP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)$
$\to\text{PC}(4\times 4,4)\to\text{Flatten}=\mathcal{F}_{\mathcal{E}}(\omega),$ (26)
where ‘PC’ stands for a convolutional layer with periodic padding and ‘MP’ for a max pooling layer. The first term in the brackets denotes the size of the convolutional filters and the second term denotes the number of filters.
The bottleneck $\mathcal{W}$ consists of a series of four fully connected (‘FC’) linear layers with equal input/output dimension,
$\mathcal{F}_{\mathcal{E}}(\omega)\to FC(1024)^{4}=\mathcal{W}(\mathcal{F}_{\mathcal{E}}(\omega)).$ (27)
The latent space is the output of this bottleneck, which is then reshaped into an image with four channels of dimension $16^{2}$. The decoder then mirrors the encoder, with the addition of an output stream of linear periodic convolutional layers (‘LPC’):
$\mathcal{W}(\mathcal{F}_{\mathcal{E}}(\omega))$ $\to\text{Reshape}$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{UP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{UP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{UP}(2\times 2)$
$\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)$
$\to\text{LPC}(16\times 16,1)\to\text{LPC}(12\times 12,1)$
$\to\text{LPC}(8\times 8,1)\to\text{LPC}(6\times 6,1)$
$=\mathcal{F}_{\mathcal{D}}(\mathcal{W}(\mathcal{F}_{\mathcal{E}}(\omega)))=\mathscr{A}(\omega),$ (28)
where ‘UP’ denotes upsampling layers.
B.2 $Re=400$
The architecture of IRMAE at $Re=400$ was slightly modified from that used at $Re=40$ and $100$, such that the input and output vorticity fields would be of resolution $512\times 512$. One extra downsampling/upsampling operation was inserted into the encoder/decoder, such that the input field was reduced to have dimension $32^{2}$, rather than $16^{2}$. The kernel sizes were also modified throughout the encoder and decoder to reflect the increased dimensionality of the input/output vorticity fields, as outlined below.
The encoder architecture is given by the following sequence of operations:
$\omega$ $\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)\to\text{MP}(2\times 2)$
$\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)\to\text{MP}(2\times 2)$
$\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)\to\text{MP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{MP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)$
$\to\text{PC}(4\times 4,1)\to\text{Flatten}=\mathcal{F}_{\mathcal{E}}(\omega),$ (29)
The input vorticity field is a single-channel image of size $512^{2}$, hence the output of the encoder $\mathcal{F}_{\mathcal{E}}(\omega)$ is a flattened vector of length $32^{2}=1024$.
The architecture of the bottleneck is unchanged from the lower $Re$ bottleneck architecture. The latent representation is then reshaped to have dimension $32^{2}$. The decoder architecture is again slightly modified to mirror the encoder,
$\mathcal{W}(\mathcal{F}_{\mathcal{E}}(\omega))$ $\to\text{Reshape}$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{UP}(2\times 2)$
$\to\text{PC}(4\times 4,32)\to\text{DB}(4\times 4)\to\text{UP}(2\times 2)$
$\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)\to\text{UP}(2\times 2)$
$\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)\to\text{UP}(2\times 2)$
$\to\text{PC}(8\times 8,32)\to\text{DB}(8\times 8)$
$\to\text{LPC}(16\times 16,1)\to\text{LPC}(12\times 12,1)$
$\to\text{LPC}(8\times 8,1)\to\text{LPC}(6\times 6,1)$
$=\mathcal{F}_{\mathcal{D}}(\mathcal{W}(\mathcal{F}_{\mathcal{E}}(\omega)))=\mathscr{A}(\omega).$ (30)
Appendix C Data Generation and IRMAE training
Datasets were generated at each $Re$ by sampling long trajectories of the flow. At $Re=\{40,100\}$, 100 trajectories of length $10^{3}$ were sampled at every advective time unit. Initial transient periods of 50 time units were discarded from the trajectories, such that the training data is sampled from the turbulent attractor. The resulting dataset was resampled according to the dissipation rate of each snapshot, as in Cleary and Page [4].
The $Re=40$ and $100$ networks were initialized with the final network weights from Cleary and Page [4], which were trained on symmetry-reduced snapshots (both continuous and discrete symmetries of Kolmogorov flow). We refer the reader to Cleary and Page [4] for full details on the training hyperparameters, symmetry reduction protocol and reconstruction error of these network weights.
However, we do not assume that the measurements are taken from symmetry-reduced turbulent states here, so these network weights are further fine-tuned on data which is augmented at every epoch with random symmetry operations. The AdamW optimizer [17] is used for this transfer learning, with an initial learning rate of $10^{-4}$ and the default weight decay coefficient of 0.004 for 500 epochs.
At $Re=400$, four trajectories of length $2.5\times 10^{4}$ were sampled at every time unit, resulting in $N_{S}=10^{5}$ snapshots. Again, initial transient periods of 50 time units were discarded from these trajectories. These snapshots were not resampled according to dissipation rate, as i) it is more computationally expensive to generate large datasets at the increased $Re$ and increased resolution of $512^{2}$ and ii) the dissipation rate distribution is less skewed at the higher $Re$.
The $Re=400$ IRMAE network was trained from scratch to input/output states on the finer computational grid of $512^{2}$. The AdamW optimizer is used with an initial learning rate of $10^{-4}$ and the default weight decay coefficient of 0.004 for 400 epochs. The training data is symmetry augmented at every epoch with random symmetry operations.
Figure 7: PDFs of the relative reconstruction error of IRMAE at $Re=40,100$ and $400$, from left to right.
The probability density functions (PDFs) of the relative errors for each snapshot $\omega_{i}$ in the test datasets at $Re=\{40,100,400\}$,
$\varepsilon_{i}=\frac{\|\omega_{i}-\mathscr{A}(\omega_{i})\|}{\|\omega_{i}\|}\,$, (31)
is reported in figure 7. Each $\omega_{i}$ in the test datasets is symmetry-augmented by random continuous and discrete symmetry operations. As expected, the reconstruction accuracy decreases with $Re$ as the spatio-temporal complexity of the flow increases. Longer tails are also clear as $Re$ increases, likely reflecting the increasing proportion of high-dissipation bursting events in the test datasets.